# Message-Passing Monte Carlo: Generating low-discrepancy point sets via Graph Neural Networks

### v1

@@ Appendix A Training details
Appendix B The Asian Option Problem Formulation
@@
In our simulations, the true value $C_{0}=7.04704$ was calculated in advance via QMC simulation with $2$M Sobol’ points with the following set of parameters: $S(0)=50,T=1\text{ year},r=0.05,\sigma=0.3,K=45$ and $d=32$.
Appendix C Further empirical insights
C.1 Structure of MPMC Points
@@
Initially explored in (Clément et al., 2023), the authors provide insights into the configurations of two dimensional point sets that achieve optimal star-discrepancy by providing visualizations of the local discrepancy within the unit square. Providing equivalent comparisons, Figure 5 displays the local discrepancy plots for Sobol’ sequences, MPMC points, and optimal point sets, respectively. Each plot has its own color scale where darker areas indicate lower local discrepancy values, and brighter areas denote higher values. The presence of a black dot in each plot marks the point of maximum local discrepancy, i.e., the explicit anchored test box where the star-discrepancy is obtained. Additionally, regions of high local discrepancy form bright triangular regions whose corner is either directed toward the upper right corner to represent open boxes with too few points, or angled toward the lower left, indicating closed boxes with an excess of points. For instance, the Sobol’ plots exclusively show closed overfilled boxes, with bright triangles pointing downward and leftward toward the origin.
@@
Figure 5: Local discrepancy plots for Sobol’ sequence (top), optimal point sets obtained in (Clément et al., 2023) (middle row), and MPMC point sets (bottom) for $N=6$ points (left), $N=12$ points (middle column), and $N=18$ points (right).
C.2 Projections of MPMC points
As noted in the main text, when applied to the computational finance integral described in Section B to estimate the value of an Asian call option, the MPMC points far outperform the Sobol’ or Halton in terms of approximation accuracy. A significantly important factor for the success of QMC methods in high dimensional application is the quality of the distribution in the lower dimensional projections of the QMC point set. See (Lemieux & Owen, 2002) for a more comprehensive discussion. Figure 6 and 7 display the $5$-th and $6$-th, and $26$-th and $27$-th coordinate projections respectively of the MPMC points, Sobol’ and Halton sequences constructed in 32 dimensions. Visual inspection reveals that the projections seem to be just as uniformly distributed in the $5$-th and $6$-th dimensions and notably more evenly distributed as the dimension increases to $26$ and $27$. At these higher dimensions, we start noticing some undesired correlations in the coordinates of the Sobol’ and Halton sequences, however, fortunately the MPMC construction does not exhibit this problematic feature displaying no significant correlation, clustering, or sparsity; the MPMC point sets appear random yet maintain a high degree of uniformity. This characteristic is particularly advantageous for tackling high-dimensional problems.
Figure 6: Projections of the $5$-th and $6$-th coordinates of 32-dimensional MPMC (left), Sobol’ (middle) and Halton (right) with $N=512$.
Figure 7: Projections of the $26$-th and $27$-th coordinates of 32-dimensional MPMC (left), Sobol’ (middle) and Halton (right) with $N=512$.
C.3 Optimality of MPMC in Three Dimensions
Table 4 shows the star-discrepancy of MPMC in three dimensions for $N=1,2,\dots,8$ number of points, as well as the optimal star-discrepancy values obtained from (Clément et al., 2023). We can see that MPMC obtains again near-optimal star-discrepancy for all choices of $N$.
@@ C.3 Optimality of MPMC in Three Dimensions
C.4 Star-discrepancy values of Figure 4 in the main text
From the main text, we present the exact numerical values of the star-discrepancy for Halton, Sobol’, Subset Selection, Hammersley, Fibonacci, and MPMC points as a benchmark for future methods.
Table 5: Star-discrepancy values of Figure 4 in the main text for Halton, Sobol’, Subset Selection, Hammersley, Fibonacci, and MPMC.
@@ C.4 Star-discrepancy values of Figure 4 in the main text
Halton 0.1738 0.0654 0.0502 0.0369 0.032 0.0232 0.0206 0.0199 0.0195 0.0166 0.0162 0.0128 0.0117
Sobol 0.1312 0.0658 0.0462 0.0331 0.0247 0.0242 0.0257 0.0185 0.0128 0.0144 0.0133 0.0123 0.0129
Subset Selection 0.0880 0.0547 0.0286 0.0244 0.0203 0.015 0.0134 0.0153 0.0117 0.0109 0.0107 0.0095 0.0090
Hammersley 0.1230 0.0494 0.0314 0.0229 0.0184 0.0151 0.0131 0.0116 0.0106 0.0094 0.0085 0.008 0.0074
Fibonacci 0.1188 0.0442 0.0275 0.0213 0.0165 0.0135 0.012 0.0105 0.0096 0.0086 0.0078 0.0071 0.0065
MPMC 0.0666 0.0273 0.0188 0.0137 0.0115 0.0097 0.0084 0.0075 0.0071 0.0070 0.0058 0.0054 0.0052
C.5 On the role of the radius in the nearest neighbor graph
@@
We recall from the main text, that our proposed MPMC method is based on GNNs that leverage $r$-radius nearest neighbors as the underlying computational graph, connecting nodes within a given radius $r$. How does the performance of MPMC depend on the radius $r$? Moreover, is it necessary to use GNNs? To answer this, we train $10$ MPMC models for different radius values $r=0.1,\dots,1.0$ for different number of points $N$ and plot the resulting average ${\mathcal{L}}_{2}$-discrepancy in Fig. 9. We can see that there is no correlation between the performance and a fixed radius $r$. Moreover, the performance appears to be not overly sensitive with respect to different values for the radius. Nevertheless, small variations of the performance with respect to the radius $r$ can be seen and it is thus advisable to include the radius to the set of tune-able hyperparameters of the model.
@@
Figure 8: ${\mathcal{L}}_{2}$-discrepancy of MPMC points for increasing values of the radius of the underlying nearest neighbor computational graph ranging from $0.1$ to $1.0$ for different number of points $N=64,128,1024$.
@@
Figure 9: ${\mathcal{L}}_{2}$-discrepancy of MPMC for different choices of GNN architectures, i.e., GCN (Kipf & Welling, 2017), GAT (Velickovic et al., 2018), and MPNN (Gilmer et al., 2017) for three different number of points $N=64,256,1024$.
@@
We further note that a radius of $r=0$ corresponds to zero edges in the underlying computational graph. Thus, the model becomes a deepset (Zaheer et al., 2017), processing each point in the set individually without aggregating any neighborhood information. The average ${\mathcal{L}}_{2}$-discrepancy of this deepset is approximately $0.073$ for $N=64$, $0.058$ for $N=128$, and $0.063$ for $N=1024$, i.e., between $9$ to over $40$ times worse than GNNs with $r\geq 0.1$. Moreover, the deepset fails to decrease the ${\mathcal{L}}_{2}$-discrepancy for increasing number of points $N$. This highlights the necessity of using GNNs that aggregate geometric information from neighboring points for successfully generating low-discrepancy point sets.
C.6 On the role of the GNN architecture
@@
While we base our proposed MPMC model on MPNNs (Gilmer et al., 2017), any other GNN architecture could be used instead. Therefore, it is natural to ask how the choice of the GNN architecture influences the performance of MPMC. To answer this, we test three different configurations of MPMC: one based on MPNNs, one based on GCNs (Kipf & Welling, 2017), and one based on GATs (Velickovic et al., 2018). We train all three configurations for different number of points $N=64,256,1024$, and provide the results as a bar plot in Fig. 9. We can see that while GCNs and GATs outperform each other depending on the chosen number of points, MPNNs consistently produce point sets with the lowest ${\mathcal{L}}_{2}$-discrepancy among all three configurations for all number of points considered here.
C.7 On the role of the input points
Figure 10: .
@@ C.7 On the role of the input points
In the main text, we suggest three different input types to be transformed into low-discrepancy points via our MPMC framework, namely random points, Sobol’, and randomized Sobol’. In this experiment, we empirically analyse how these different types influence the discrepancy of the resulting MPMC points. To this end, we train several MPMC models based on the three different input types and report the average ${\mathcal{L}}_{2}$-discrepancy during training for two different number of points $N=256,1024$ in Fig. 10. We can see that on average either Sobol’ or randomized Sobol’ reach lower discrepancy values as well as exhibit faster convergence compared to random points. We note, however, that the single best MPMC model for each of the three different input point types yield almost identical discrepancy values. Thus, we conclude that Sobol’ and randomized Sobol’ points on average yield lower discrepancy values compared to random points, while at the same time the best performing input type has to be evaluated in practice for each number of points $N$.

### v2

@@ Appendix A Training details
Appendix B On the Asian Option Pricing Experiment
@@
In our simulations, the true value $C_{0}=7.06574$ was calculated in advance via QMC simulation with $2$M Sobol’ points with the following set of parameters: $S(0)=50,T=1\text{ year},r=0.05,\sigma=0.3,K=45$ and $d=32$.
Further, for each $N$, the generating vector for the rank-1 lattice is produced from LatNet builder from the command line by the following syntax: latnetbuilder -t lattice -c ordinary -s N -d 32 -e full-CBC -f CU:P2 -q 2 --weights file:/path/to/weights.txt. The file weights.txt contains order-dependent weights of $10$ on projection orders 1, 2 and 3 and otherwise a default weight of $0.001$.
Appendix C On the $\mathcal{L}_{2}$-discrepancy
The ${\mathcal{L}}_{2}$-discrepancy is a well-researched and widely used measure of distribution irregularity, consistently attracting attention from the QMC research community (63, 64, 65). In our MPNN framework, we utilize the $\mathcal{L}_{2}$-discrepancy function as a training objective because of its efficient computation and differentiable formulation to be used in the gradient-based learning. This allows the generation of point sets with very small star discrepancy. As a by-product, we also create point sets with small $\mathcal{L}_{2}$-discrepancy. Fig. 5 shows the ${\mathcal{L}}_{2}$-discrepancy values for increasing number of points $N=20,\dots,1020$ for MPMC, Halton, Sobol’, Subset Selection method, Hammersley, and Fibonacci in $2$ dimensions. We further present the numerical values of Fig. 5 in Table 4. We can see that MPMC consistently obtains the lowest ${\mathcal{L}}_{2}$-discrepancy values for all number of points $N$. Moreover, MPMC significantly outperforms the previous state-of-the-art method based on Fibonacci point sets. We highlight that the performance gap between MPMC and any other method considered here is even wider for the ${\mathcal{L}}_{2}$-discrepancy than for the star-discrepancy $D^{*}$ in Figure 4 of the main text. This can be explained by the fact that MPMC is trained to minimize the ${\mathcal{L}}_{2}$-discrepancy directly, and thus by design optimizes the ${\mathcal{L}}_{2}$-discrepancy instead of the star-discrepancy $D^{*}$.
Table 4: $\mathcal{L}_{2}$-discrepancy values for Halton, Sobol’, Subset Selection, Hammersley, Fibonacci, and MPMC for different number of points $N=20,\dots,1020$ in $d=2$.
N 20 60 100 140 180 220 260 300 340 380 420 460 500
Halton 0.06511 0.01323 0.00751 0.00627 0.00460 0.00434 0.00389 0.00343 0.00292 0.00238 0.00266 0.00215 0.00226
Sobol’ 0.03564 0.01660 0.01006 0.00645 0.00551 0.00690 0.00655 0.00435 0.00276 0.00330 0.00267 0.00258 0.00242
Subset Selection 0.02569 0.01541 0.00823 0.00693 0.00530 0.00397 0.00352 0.00410 0.00310 0.00276 0.00266 0.00250 0.00221
Hammersley 0.04796 0.01812 0.01119 0.00833 0.00664 0.00554 0.00469 0.00416 0.00371 0.00335 0.00305 0.00283 0.00262
Fibonacci 0.04324 0.01465 0.00870 0.00657 0.00492 0.00399 0.00344 0.00306 0.00275 0.00249 0.00221 0.00198 0.00182
MPMC 0.02016 0.00756 0.00479 0.00353 0.00284 0.00241 0.00203 0.00179 0.00162 0.00154 0.00135 0.00122 0.00117
N 540 580 620 660 700 740 780 820 860 900 940 980 1020
Halton 0.00182 0.00179 0.00164 0.00171 0.00190 0.00179 0.00139 0.00122 0.00139 0.00126 0.00147 0.00117 0.00107
Sobol’ 0.00220 0.00216 0.00157 0.00152 0.00198 0.00164 0.00138 0.00196 0.00155 0.00154 0.00130 0.00125 0.00121
Subset Selection 0.00206 0.00228 0.00205 0.00178 0.00166 0.00167 0.00141 0.00151 0.00153 0.00152 0.00136 0.00112 0.00120
Hammersley 0.00243 0.00224 0.00211 0.00200 0.00190 0.00181 0.00173 0.00165 0.00158 0.00152 0.00146 0.00140 0.00135
Fibonacci 0.00168 0.00155 0.00145 0.00139 0.00132 0.00127 0.00121 0.00115 0.00110 0.00107 0.00103 0.00099 0.00094
MPMC 0.00104 0.00104 0.00098 0.00090 0.00087 0.00088 0.00082 0.00080 0.00074 0.00075 0.00072 0.00072 0.00068
Figure 5: $\mathcal{L}_{2}$-discrepancy of Halton, Sobol’, Subset Selection, Hammersley, Fibonacci, and MPMC for increasing number of points $N=20,\dots,1020$ in $d=2$.
Appendix D Further empirical insights
D.1 Post-construction enhancements
Classical quasi-Monte Carlo (QMC) constructions have known limitations, including poor uniformity in low-dimensional projections of high-dimensional point sets (see Figure 7 and 8). To mitigate these issues, significant amount of research has concentrated on incorporating randomness into these deterministic constructions, as discussed in (66, 67, 68). This approach also offers the advantage of using repeated uniform sampling – known as randomized quasi-Monte Carlo (RQMC) – for straightforward unbiased error estimation.
One successful method is Owen scrambling in base $b$ (69) which involves sequentially applying uniformly chosen permutations to the base $b$ digits of each point in the set. This method has proven to be very popular and highly effective. Another simpler technique is the random shift modulo $1$ (70), where a vector in $[0,1]^{d}$ is chosen uniformly at random and added (modulo 1) to each point in the set.
In this section, we test how these enhancements improve the results of selected QMC constructions in the Asian option pricing application presented in Section 3.C. of the main text (i.e., Table 2 of the main text). To this end, we start by applying scrambling (in the natural base $2$) to the Sobol’ sequence and present the mean absolute error (MAE) in Table 5. We can see that on average scrambling massively improves the absolute error of Sobol’, i.e., by a factor of more than $10$. While Owen scrambling in base $b$ can be applied to an arbitrary point set, the choice of $b$ is paramount to ensure that the low-discrepancy property is preserved after randomization. Since the correct choice of the base is not immediately apparent, scrambling is not directly applicable to MPMC and a direct comparison involving both methods leveraging the scrambling enhancement is not possible. However, uniform random shifting (modulo 1) can be applied to any QMC construction, including MPMC. Therefore, we present the results of randomly shifted MPMC in Table 5. We can see that randomly shifting on average leads to a lower absolute error compared to MPMC without random shifting for smaller number of points $N$. For large $N=1024$, however, MPMC with random shifting appears to perform worse than MPMC without random shifting. This highlights the importance of developing suitable randomization techniques specifically tailored for MPMC, a topic we plan to focus on in future research.
Table 5: Errors of Sobol’, scrambled Sobol’, MPMC and randomly shifted MPMC for the Asian option pricing experiment in Section 3.C. of the main text. MPMC and Sobol’ errors are provided as absolute errors and taken from Table 2 in the main text, while shifted MPMC and scrambled Sobol’ are provided as mean absolute errors (MAE).
N 32 64 128 256 512 1024
Sobol’ 1.235 1.373 0.965 0.623 0.497 0.290
Scrambled Sobol’ 0.516 0.169 0.076 0.064 0.040 0.022
MPMC 1.402 0.831 0.512 0.250 0.120 0.055
Shifted MPMC 0.521 0.310 0.188 0.128 0.082 0.061
Figure 6: Local discrepancy plots for Sobol’ sequence (top), optimal point sets obtained in (29) (middle row), and MPMC point sets (bottom) for $N=6$ points (left), $N=12$ points (middle column), and $N=18$ points (right).
D.2 Structure of MPMC Points
@@
Initially explored in (29), the authors provide insights into the configurations of two dimensional point sets that achieve optimal star-discrepancy by providing visualizations of the local discrepancy within the unit square. Providing equivalent comparisons, Figure 6 displays the local discrepancy plots for Sobol’ sequences, optimal point sets and MPMC points. Each plot has its own color scale where darker areas indicate lower local discrepancy values, and brighter areas denote higher values. The presence of a black dot in each plot marks the point of maximum local discrepancy, i.e., the explicit anchored test box where the star-discrepancy is obtained. Additionally, regions of high local discrepancy form bright triangular regions whose corner is either directed toward the upper right corner to represent open boxes with too few points, or angled toward the lower left, indicating closed boxes with an excess of points. For instance, the Sobol’ plots exclusively show closed overfilled boxes, with bright triangles pointing downward and leftward toward the origin.
@@
D.3 Projections of MPMC points
As noted in the main text, when applied to the computational finance integral described in Section B to estimate the value of an Asian call option, the MPMC points far outperform the Sobol’ or Hammersley in terms of approximation accuracy. A significantly important factor for the success of QMC methods in high dimensional application is the quality of the distribution in the lower dimensional projections of the QMC point set. See (52) for a more comprehensive discussion. Figure 7 and 8 display the $5$-th and $6$-th, and $26$-th and $27$-th coordinate projections respectively of the MPMC points, Sobol’, rank-1 lattice and Hammersley constructed in $32$ dimensions. Visual inspection reveals that the projections seem to be just as uniformly distributed in the $5$-th and $6$-th dimensions and notably more evenly distributed as the dimension increases to $26$ and $27$. At these higher dimensions, we start noticing some undesired correlations in the coordinates of the Sobol’ and Hammersley constructions, however, fortunately the MPMC construction does not exhibit this problematic feature displaying no significant correlation, clustering, or sparsity; the MPMC point sets appear random yet maintain a high degree of uniformity. This characteristic is particularly advantageous for tackling high-dimensional problems.
Figure 7: Projections of the $5$-th and $6$-th coordinates of 32-dimensional MPMC, Sobol’, rank-1 lattice and Hammersley with $N=512$ (left to right).
Figure 8: Projections of the $26$-th and $27$-th coordinates of 32-dimensional MPMC, Sobol’, rank-1 lattice and Hammersley with $N=512$ (left to right).
D.4 Optimality of MPMC in Three Dimensions
Table 7 shows the star-discrepancy of MPMC in three dimensions for $N=1,2,\dots,8$ number of points, as well as the optimal star-discrepancy values obtained from (29). We can see that MPMC obtains again near-optimal star-discrepancy for all choices of $N$.
@@ D.4 Optimality of MPMC in Three Dimensions
D.5 Star-discrepancy values of Figure 4 in the main text
From the main text, we present the exact numerical values of the star-discrepancy for Halton, Sobol’, Subset Selection, Hammersley, lifted Sobol’, Fibonacci, and MPMC points as a benchmark for future methods.
Table 7: Star-discrepancy values of Figure 4 in the main text for Halton, Sobol’, Subset Selection, Hammersley, lifted Sobol’, Fibonacci, and MPMC.
@@ D.5 Star-discrepancy values of Figure 4 in the main text
Halton 0.17384 0.06535 0.05024 0.03686 0.03200 0.02323 0.02062 0.01994 0.01950 0.01659 0.01617 0.01279 0.01172
Sobol’ 0.13125 0.06583 0.04617 0.03306 0.02466 0.02420 0.02571 0.01851 0.01280 0.01442 0.01326 0.01225 0.01289
Subset Selection 0.08799 0.05469 0.02860 0.02439 0.02028 0.01499 0.01339 0.01527 0.01175 0.01089 0.01073 0.00950 0.00898
Hammersley 0.12304 0.04941 0.03136 0.02292 0.01842 0.01512 0.01308 0.01164 0.01056 0.00945 0.00855 0.00803 0.00739
Lifted Sobol’ 0.11875 0.04609 0.02766 0.02388 0.01501 0.01584 0.01221 0.01294 0.00837 0.00994 0.00915 0.00849 0.00721
Fibonacci 0.11885 0.04422 0.02749 0.02128 0.01655 0.01354 0.01200 0.01054 0.00957 0.00857 0.00775 0.00708 0.00651
MPMC 0.06664 0.02729 0.01879 0.01373 0.01147 0.00975 0.00843 0.00752 0.00710 0.00695 0.00584 0.00540 0.00518
N 540 580 620 660 700 740 780 820 860 900 940 980 1020
Halton 0.01101 0.01005 0.00957 0.00949 0.00841 0.00709 0.00685 0.00718 0.00710 0.00571 0.00626 0.00619 0.00636
Sobol’ 0.00967 0.00967 0.00977 0.01064 0.00724 0.00765 0.00697 0.00691 0.00772 0.00638 0.00637 0.00640 0.00616
Subset Selection 0.00800 0.00927 0.00802 0.00727 0.00680 0.00664 0.00582 0.00626 0.00626 0.00602 0.00536 0.00457 0.00488
Hammersley 0.00685 0.00637 0.00596 0.00578 0.00545 0.00523 0.00496 0.00472 0.00450 0.00431 0.00413 0.00396 0.00380
Lifted Sobol’ 0.00642 0.00645 0.00534 0.00759 0.00509 0.00466 0.00547 0.00441 0.00422 0.00450 0.00373 0.00471 0.00362
Fibonacci 0.00603 0.00561 0.00525 0.00509 0.00480 0.00484 0.00459 0.00436 0.00416 0.00398 0.00381 0.00365 0.00351
MPMC 0.00488 0.00476 0.00454 0.00437 0.00416 0.00397 0.00376 0.00360 0.00356 0.00319 0.00321 0.00308 0.00303
D.6 On the role of the radius in the nearest neighbor graph
@@
We recall from the main text, that our proposed MPMC method is based on GNNs that leverage $r$-radius nearest neighbors as the underlying computational graph, connecting nodes within a given radius $r$. How does the performance of MPMC depend on the radius $r$? Moreover, is it necessary to use GNNs? To answer this, we train $10$ MPMC models for different radius values $r=0.1,\dots,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sqrt{2}}$ for different number of points $N$ in $d=2$ and plot the resulting average ${\mathcal{L}}_{2}$-discrepancy in Fig. 10. We can see that there is no correlation between the performance and a fixed radius $r$. Moreover, the performance appears to be not overly sensitive with respect to different values for the radius. Nevertheless, small variations of the performance with respect to the radius $r$ can be seen and it is thus advisable to include the radius to the set of tune-able hyperparameters of the model.
@@
Figure 9: ${\mathcal{L}}_{2}$-discrepancy of MPMC points for increasing values of the radius of the underlying nearest neighbor computational graph ranging from $0.1$ to ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sqrt{2}}$ for different number of points $N=64,128,1024$ in $d=2$.
@@
Figure 10: ${\mathcal{L}}_{2}$-discrepancy of MPMC for different choices of GNN architectures, i.e., GCN (38), GAT (71), and MPNN (42) for three different number of points $N=64,256,1024$ in $d=2$.
@@
We further note that a radius of $r=0$ corresponds to zero edges in the underlying computational graph. Thus, the model becomes a deepset (62), processing each point in the set individually without aggregating any neighborhood information. The average ${\mathcal{L}}_{2}$-discrepancy of this deepset is approximately $0.073$ for $N=64$, $0.058$ for $N=128$, and $0.063$ for $N=1024$, i.e., between $9$ to over $40$ times worse than GNNs with $r\geq 0.1$. Moreover, the deepset fails to decrease the ${\mathcal{L}}_{2}$-discrepancy for increasing number of points $N$. This highlights the necessity of using GNNs that aggregate geometric information from neighboring points for successfully generating low-discrepancy point sets.
D.7 On the role of the GNN architecture
@@
While we base our proposed MPMC model on MPNNs (42), any other GNN architecture could be used instead. Therefore, it is natural to ask how the choice of the GNN architecture influences the performance of MPMC. To answer this, we test three different configurations of MPMC: one based on MPNNs, one based on GCNs (38), and one based on GATs (71). We train all three configurations for different number of points $N=64,256,1024$ in $d=2$, and provide the results as a bar plot in Fig. 10. We can see that while GCNs and GATs outperform each other depending on the chosen number of points, MPNNs consistently produce point sets with the lowest ${\mathcal{L}}_{2}$-discrepancy among all three configurations for all number of points considered here.
Figure 11: ${\mathcal{L}}_{2}$-discrepancy values of MPMC during training for three different types of input points, i.e., Sobol’, Randomized Sobol’, and random points, for $N=256$ and $N=1024$ in $d=2$.
D.8 On the role of the input points
@@
In the main text, we suggest three different input types to be transformed into low-discrepancy points via our MPMC framework, namely random points, Sobol’, and randomized Sobol’. In this experiment, we empirically analyse how these different types influence the discrepancy of the resulting MPMC points. To this end, we train several MPMC models in $d=2$ based on the three different input types and report the average ${\mathcal{L}}_{2}$-discrepancy during training for two different number of points $N=256,1024$ in Fig. 11. We can see that on average either Sobol’ or randomized Sobol’ reach lower discrepancy values as well as exhibit faster convergence compared to random points. We note, however, that the single best MPMC model for each of the three different input point types yield almost identical discrepancy values. Thus, we conclude that Sobol’ and randomized Sobol’ points on average yield lower discrepancy values compared to random points, while at the same time the best performing input type has to be evaluated in practice for each number of points $N$.